[Brahms-l] Shift report 20050324 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Thu Mar 24 2005 - 08:02:03 EST
Shifter leader was JHL
Shifters were 


- The Last shift for Cu-Cu.

- Beam at the injection energy for 6h40m.
  Store lengh: ~1h15m
  Time between stores: 8m-30m

- No detector Problems.

==============================================
=====Summary of the Injection Energy Run======
==============================================
- Beam: Cu29+ on Cu29+ at 11.2 GeV/n
- ZDC Rates: 150 ~ 10/sec

- Triggers Used:
Trigger1 BB (>=1)
Trigger2 FS (with RC)
Trigger3 MRS (with RC)
Trigger4 ZDC (at the same HV settings as 62 GeV run)
Trigger5 RC+Tile>=1
Trigger6 FFS

- Toral DAQ time: ~26.5 hours (Mar. 22 - Mar. 24)

- Data Runs 13928 - 13993
MRS 40deg 700B: 1330K     Trigger3
    45deg 700B:  450K     Trigger3
    90deg 500A:   67K     Trigger3
FS  8deg 430B:  33K/187K Trigger2/Trigger6
    4deg 861B: 170K/668K Trigger2/Trigger6
    4deg 1219B: 85K/359K Trigger2/Trigger6
    8deg 430A:  10K/78K  Trigger2/Trigger6
    4deg 861A:  17K/315K Trigger2/Trigger6
    4deg 1219A: 10K/239K Trigger2/Trigger6

===============================================



_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Mar 24 08:02:17 2005

This archive was generated by hypermail 2.1.8 : Thu Mar 24 2005 - 08:02:24 EST